OpenStudy (samigupta8):

If a,b,c are positive real number such that b+c-a,c+a-b and a+b-c are positive the expression (b+c-a)(c+a-b)(a+b-c) is A).positive B).negative C).non-positive D).nonnegative

OpenStudy (samigupta8):

Oh wait it's wrong question..

OpenStudy (samigupta8):

-abc is also there in the question

ILovePuppiesLol (ilovepuppieslol):

where will "-abc" come? please edit your question

OpenStudy (samigupta8):

(b+c-a)(c+a-b)(a+b-c)-abc

ILovePuppiesLol (ilovepuppieslol):

we can simply assume appropriate values for a,b and c and then get the answer..

OpenStudy (samigupta8):

I want a proper ans though..

OpenStudy (samigupta8):

No approximation pls

ILovePuppiesLol (ilovepuppieslol):

thats not approximation because there is only 1 solution to this question which must hold true for all values of a,b,c which follow those 3 equations.. you can do this too-assume a,b,c to be side lengths of a triangle

ILovePuppiesLol (ilovepuppieslol):

by 3 equations i mean a+b-c>0 b+c-a>0 c+a-b>0 triangle inequalities

OpenStudy (samigupta8):

Bt afterwards we have to Assume some value of a,b,c which satisfy these three equation..

OpenStudy (samigupta8):

@ilovepuppieslol do u know of another way??

ILovePuppiesLol (ilovepuppieslol):

what are your thoughts on this question? :|

OpenStudy (samigupta8):

Lol...

ILovePuppiesLol (ilovepuppieslol):

...

OpenStudy (samigupta8):

I opened up the brackets ..and came up with a very big expression

ILovePuppiesLol (ilovepuppieslol):

\(a^3 +b^3+c^3-3abc=(a + b + c)(a^2 + b^2 + c^2 - ab - bc - ac) \) \(\Large \frac{a^3 +b^3+c^3-(a + b + c)(a^2 + b^2 + c^2 - ab - bc - ac) }{3}=abc\) \((b+c-a)(c+a-b)(a+b-c)-abc\) \((b+c-a)(c+a-b)(a+b-c)-\Large\frac{a^3 +b^3+c^3-(a + b + c)(a^2 + b^2 + c^2 - ab - bc - ac) }{3}\)

OpenStudy (samigupta8):

Did u do it like that?

ILovePuppiesLol (ilovepuppieslol):

no i just shared another form of the equation

OpenStudy (samigupta8):

Oh .i see

OpenStudy (samigupta8):

Well. I feel now that your first approach to the problem was nice... For if i solve it like this it wud be very time consuming...

ILovePuppiesLol (ilovepuppieslol):

yeah a typical way of doing this would involve expanding the stuff and that would take time :(

OpenStudy (samigupta8):

Thanks

ILovePuppiesLol (ilovepuppieslol):

np